OpenStudy (anonymous):

How do I use integration by parts to integrate this problem (p^3)(e^(p^2)) from 0 to 1

OpenStudy (amistre64):

rule of thumb, a poly always drives to zero in the end ...

OpenStudy (amistre64):

**derives to zero that is

OpenStudy (amistre64):

@zepdrix i was curious to see what you were responding with, which is why i didnt make any indepth analysises. just a commentary.

OpenStudy (anonymous):

So I used the poly as u and I got (p^2*e^(p^2))/2 - 3/2 * (integral of p*e^(p^2))) @amistre64

OpenStudy (anonymous):

Sorry about the late reply

OpenStudy (amistre64):

the sequence that is generated can be formulated in a table, the derivatives verses the integrations sign du v 2/2 p e^(p^2) + p^2 e^(p^2) / 2 - 2p i think this might be a better view of whats going on

OpenStudy (amistre64):

\[\frac12\int \underbrace{p^2}_{u}~2\underbrace{p~e^{p^2}dp}_{dv}=\frac12\left[p^2~e^{p^2}-\int2 p~e^{p^2}dp \right]\]

OpenStudy (anonymous):

gotcha thanx for clearing that up :)